# TiTok: Transfer Token-level Knowledge via Contrastive Excess to Transplant LoRA

### v1

@@ 4.2 Main results
Table 1: Main results. Experiments on BBH, MMLU, News Headline and Scholarly Title Generation tasks under four transfer settings. BBH and MMLU are reasoning tasks and are evaluated using LM-Eval Harness, while News Headline and Scholarly Title Generation represent personalization tasks and are evaluated with ROUGE-1/L. All evaluations are zero-shot. Best scores are in bold.
@@
Table 1 summarizes the experimental results on BBH, MMLU, and LaMP (News Headline and Scholarly Title generation) across four transfer settings. First, we observe that transfer within the same model family is highly effective; when transplanting the LoRA adapter trained on the Mistral 7B into a fresh instance of the same model, TiTok consistently surpasses all baselines. In particular, TiTok improves the vanilla model by 24.08% on average across all tasks, and further outperforms the KD and TransLoRA baselines by 19.63% and 4.91%, respectively. Taken together, the findings show that, as a first step, transfer within the same family is reliably successful.
Beyond same family transfer, we also find that TiTok is highly effective in cross-model transfer settings. For example, when transferring from Mistral 7B to Llama 8B, TiTok delivers an average improvement across all tasks of 7.11% over the vanilla model, while outperforming KD by 4.73% and TransLoRA by 6.24%. This highlights that TiTok is not confined to intra-family knowledge transfer, but can successfully bridge across architectures. In the case of Llama3 3B to Llama3 8B, TiTok yields average gains of 3.45% against vanilla, 2.49% against KD, and 4.14% against TransLoRA. These results suggest that TiTok scales effectively with model size, making it useful when moving from lightweight to larger models without losing efficiency. Finally, when transferring from Llama2 7B to Llama3 8B, TiTok performs average advantages of 7.43% over vanilla, 6.28% over KD, and 7.22% over TransLoRA. This indicates that TiTok adapts robustly even across different model versions, implying practical relevance when models are upgraded in real-world pipelines.
@@
Table 2: Ablation study. ”Sample filtering” uses the mean token-level contrastive excess score to remove uninformative examples, while “Token selection” further refines the retained data by selectively keeping only the most informative tokens within each sample. Without sample filtering, data are randomly sampled. Results are reported as accuracy (Acc.) for BBH and MMLU, averaged across tasks. Meanwhile, we report ROUGE-1 (R-1) and ROUGE-L (R-L) for LaMP tasks.
@@
4.3 Additional analyses
Ablation study.
To validate the contribution of each component in our framework, we conduct additional experiments by selectively excluding the sample filtering and token selection mechanisms in Sec. 3.3. We report the average performance scores across all four transfer settings, and the results are presented in Table 2. When sample filtering is not applied (1st and 2nd rows), the training data for the target model is randomly sampled from the full synthetic dataset. Under this setting, applying only token selection (2nd row) performs noticeably better than the purely random baseline. This demonstrates the effectiveness of our token selection approach and empirically confirms that contrastive excess reliably identifies and selects the most informative tokens. Similarly, when only sample filtering is applied (1st and 3rd rows), the results improve over pure random sampling, indicating that selecting high-quality data is essential. This finding further underscores that incorporating an effective filtering mechanism for synthetic data is essential. In our framework, contrastive excess serves this role by reliably selecting samples with richer and more informative signals, as demonstrated clearly by the empirical results Finally, the results show that combining both stages (4th row) achieves the best performance, confirming their complementary roles in filtering high-quality examples and selecting informative tokens for effective knowledge transfer.
Impact of query generation model.
We now move on to examine how the choice of query generation model influences performance. Practically, synthetic queries can be generated either by the source model or by the target model, with the latter being the approach originally adopted in TransLoRA’s pipeline (Wang et al., 2024). Figure 2 compares these two options across different transfer settings, with the reported scores averaged over all BBH tasks. Notably, we observe that using the source expert model (source backbone + LoRA) for query generation generally yields stronger performance than using the target model. One possible explanation is that when both the synthetic queries and the corresponding labels are generated by the same model, they remain closer to its training distribution. This alignment between queries and labels likely makes the supervision more coherent and accurate, thereby facilitating more effective transfer. These results suggest that maintaining distributional alignment between queries and labels is a key factor for improving knowledge transfer, thereby providing empirical justification for our design choice of using the source model as both the synthetic query and label generator.
@@ 4.3 Additional analyses
We now proceed to explore the impact of token selection ratios ($k\%$) in our framework. Fig. 3 presents three representative curves: one task (BBH) under two transfer settings, and another task (News Headline Generation) under the same transfer setting for comparison. For BBH with the same backbone on source and target (Mistral 7B→Mistral 7B), a moderate token selection ratio of 40-70% yields the best results. Very low ratios (10-20%) lead to underfitting, while a 100% ratio is ineffective as it applies no filtering at all. Interestingly, in a weak-to-strong transfer (Llama2 7B → Llama3 8B) for the same BBH task, performance improves as k% decreases. This suggests that selecting only the tokens with the highest excess loss effectively filters out the majority of noisy regions where the weaker model is uncertain. By discarding the majority of tokens where the weaker model lacks confidence, TiTok can significantly reduce negative transfer. This trend, however, reverses for the News Headline Generation task under the same weak-to-strong transfer setting. For this task, a larger k% is generally better, potentially because even a weaker model can provide valuable lexical and stylistic cues that are useful for personalization. Consequently, unlike reasoning-focused tasks such as BBH, which are sensitive to noisy supervision, personalization tasks like News Headline Generation benefit more broadly from the source model’s outputs, even when the source model is relatively weaker than the target model.
@@
Table 3: Transfer using external data. ROUGE-1 (R-1) and ROUGE-L (R-L) on News Headline and Scholarly Title Generation under three data settings for Mistral 7B→Mistral 7B: (1) other-user, (2) mixed-user, (3) cross-task. Best scores are in bold.
@@ 5 Conclusion
In this paper, we propose TiTok, an efficient framework that transfers LoRA knowledge from a source model to a target model by training only on a selectively chosen set of highly informative tokens. At its core, TiTok leverages token-level signals to distill task-relevant information from the source adapter, rather than relying on the entire token sequence. By focusing supervision on the regions where the adapter contributes most, TiTok achieves stronger and more targeted knowledge transfer. With this simple yet effective design, TiTok proves robust across tasks and consistently surpasses existing baselines, making it a practical solution for efficient knowledge transfer.
Limitations and Future Directions
While TiTok has shown robustness and clear advantages over existing methods, there remain opportunities to refine and extend the framework. Firstly, although TiTok is designed to minimize dependence on original data, it still requires a small number of seed examples to generate synthetic data via prompting. Nevertheless, this reliance is modest, as we avoid using full datasets. In addition, our experiments confirm that external data can also serve as an effective alternative, reinforcing the flexibility of the approach. Regarding token selection, TiTok currently applies a fixed threshold to determine which tokens are retained. While this simple design is effective and stable across tasks, future work could explore more adaptive or data-driven thresholding strategies to further enhance efficiency without compromising robustness.
@@
We provide a comprehensive description of our implementation in Section 4, including pipeline configurations, hyperparameters, models, datasets, and evaluation metrics. The source code for our implementation and experiments will be made publicly available in a repository upon publication.
@@ Appendix A Tokenizer Alignment Algorithm
This process yields fractional mask scores that capture the relative importance of each target token. Finally, we keep only the top-$k\%$ of tokens according to these scores, producing a final binary selection mask over the target tokens that preserves the most informative regions while discarding less relevant ones. An overview of this tokenizer alignnment algorithm is presented in Figure 4.
Appendix B Details of Datasets
B.1 Big-Bench Hard (BBH)
@@
STEM (29) abstract_algebra, anatomy, astronomy, college_biology, college_chemistry, college_computer_science, college_mathematics, college_medicine, college_physics, computer_security, conceptual_physics, electrical_engineering, elementary_mathematics, formal_logic, high_school_biology, high_school_chemistry, high_school_computer_science, high_school_mathematics, high_school_physics, high_school_statistics, machine_learning, medical_genetics, nutrition, professional_medicine, professional_psychology, virology, clinical_knowledge, human_aging, human_sexuality
Humanities (16) business_ethics, formal_fallacies, jurisprudence, logical_fallacies, philosophy, prehistory, world_religions, moral_disputes, moral_scenarios, professional_law, professional_accounting, high_school_world_history, high_school_us_history, high_school_european_history, global_facts, security_studies
Social Sciences (10) econometrics, high_school_macroeconomics, high_school_microeconomics, management, marketing, public_relations, sociology, us_foreign_policy, high_school_geography, high_school_government_and_politics
Other (2) miscellaneous, global_facts
B.2 Massive Multitask Language Understanding (MMLU)
Massive Multitask Language Understanding (MMLU) (Hendrycks et al., 2021) is a comprehensive benchmark for evaluating model performance across a broad range of knowledge intensive tasks. The benchmark consists of multiple choice questions and, similarly to BBH, we also apply a 90%/10% split of the original test-only data. All the 57 subtasks are categorized in Table 5.
B.3 LaMP tasks
@@
  • News Headline Generation (LaMP 4). Given an author profile consisting of previously written headlines, the model is asked to generate a headline for a new news article. The task evaluates if the model can adapt its output to reflect the author’s characteristic style in journalistic writing.
@@
Appendix C Baseline Details
@@ C.1 Vanilla
C.2 Knowledge Distillation (KD) (+MinED)
@@
C.3 TransLoRA
The TransLoRA baseline (Wang et al., 2024) transfers LoRA knowledge by generating synthetic data. In this method, the vanilla target model synthesizes queries, and the source model with its LoRA adapter provides the corresponding labels. Subsequently, a discriminator, separately trained with the source model as its base, is applied to filter the synthetic data used for fine-tuning the target model’s LoRA adapter. For the sake of consistency and fair comparison,we adopt the same hyperparameter settings from our experiments when applying the TransLoRA baseline.
Appendix D Usage of AI Assistants
AI assistants are minimally used in this work, restricted solely to language refinement. such as grammar correction, punctuation, and sentence structure. All research ideas, methodologies, and analyses are the original contributions of the authors. Thus, the use of AI is confined to editorial support and did not influence the originality or intellectual contributions of the work.

### v2

@@ 4.2 Main results
Table 1: Main results. Experiments on BBH, MMLU, News Headline and Scholarly Title Generation tasks under four transfer settings. Reasoning tasks (BBH, MMLU) are evaluated via LM-Eval Harness, and personalization tasks (News Headline, Scholarly Title Generation) via ROUGE-1/L. Zero-shot evaluations are reported as the mean across 3 random seeds. Best scores are in bold.
@@
Table 1 summarizes the experimental results on BBH, MMLU, and LaMP (News Headline and Scholarly Title generation) tasks across four transfer settings. First, we observe that transfer within the same model family is highly effective; when transplanting the LoRA adapter trained on the Mistral 7B into a fresh instance of the same model, TiTok consistently surpasses all baselines. In particular, TiTok improves the vanilla model by 23.94% on average across all tasks, and further outperforms the KD and TransLoRA baselines by 21.95% and 4.75%, respectively. Taken together, the findings show that, as a first step, transfer within the same family is reliably successful.
Beyond same family transfer, we also find that TiTok is highly effective in cross-model transfer settings. For example, when transferring from Mistral 7B to Llama 8B, TiTok delivers an average improvement across all tasks of 6.79% over the vanilla model, while outperforming KD by 4.69% and TransLoRA by 4.86%. This highlights that TiTok is not confined to intra-family knowledge transfer, but can successfully bridge across architectures. In the case of Llama3 3B to Llama3 8B, TiTok yields average gains of 3.07% against vanilla, 2.18% against KD, and 3.02% against TransLoRA. These results suggest that TiTok scales effectively with model size, making it useful when moving from lightweight to larger models without losing efficiency. Finally, when transferring from Llama2 7B to Llama3 8B, TiTok performs average advantages of 5.95% over vanilla, 5.17% over KD, and 5.13% over TransLoRA. This indicates that TiTok adapts robustly even across different model versions, implying practical relevance when models are upgraded in real-world pipelines.
@@
4.3 Additional analyses
In the following sections, we present additional analyses to showcase the effectiveness of TiTok. Unlike the main table which reports the average of three random seeds to ensure statistical stability, the empirical results discussed in the subsequent sections are reported using a single, fixed seed.
Ablation study.
To validate the contribution of each component in our framework, we conduct additional experiments by selectively excluding the sample filtering and token selection mechanisms in Sec. 3.3. We report the average performance scores across all four transfer settings, and the results are presented in Table 2. When sample filtering is not applied (1st and 2nd rows), the training data for the target model is randomly sampled from the full synthetic dataset. Under this setting, applying only token selection (2nd row) performs noticeably better than the purely random baseline. This demonstrates the effectiveness of our token selection approach and empirically confirms that token-wise contrastive excess reliably identifies and selects the most informative tokens. Similarly, when only sample filtering is applied (1st and 3rd rows), the results improve over pure random sampling, indicating that selecting high-quality data is essential. This finding further implies that incorporating an effective filtering mechanism for synthetic data is essential. In our framework, token-wise contrastive excess serves this role by reliably selecting samples with richer and more informative signals, as demonstrated clearly by the empirical results. Finally, the results show that combining both stages (4th row) achieves the best performance, confirming their complementary roles in filtering high-quality examples and selecting informative tokens for effective knowledge transfer.
Table 2: Ablation study. ”Sample filtering” uses the mean token-wise contrastive excess score to remove uninformative examples, while “Token selection” further refines the retained data by selectively keeping only the most informative tokens within each sample. Without sample filtering, data are randomly sampled. Results are reported as accuracy (Acc.) for BBH and MMLU, averaged across tasks. Meanwhile, we report ROUGE-1 (R-1) and ROUGE-L (R-L) for LaMP tasks.
@@ 4.3 Additional analyses
Next, we also evaluate whether high excess-score tokens indeed carry concentrated transferable knowledge. We verify this empirically by contrasting token subgroups of the top 20%, bottom 20%, and random 20% on BBH in the Mistral 7B → Llama3 8B transfer setting. To prevent unintentional overlap across groups and to extract the most informative region, we employ a comparatively low 20% threshold for this analysis. The results are presented in Table 3. Analytically, the empirical results show that the top 20% produces the best transfer performance, validating tat tokens with high token-wise contrastive excess scores indeed contain concentrated task knowledge.
Table 3: Excess-score token subgroup analysis. Effect of token subgroup selection in Mistral 7B $\rightarrow$ Llama3 8B transfer. The top 20% excess-score tokens yield the best performance on BBH.
[…]
Interestingly, we observe that truly important tokens are consistently selected regardless of the expert’s model. For example, for BBH datasets, Mistral 7B and Llama2 7B experts agree on 59.76% of the chosen tokens even when we limit the selection to a small top 20% subset to prevent trivial overlap. This implies that a token will continue to be recognized across models if it is truly significant for the task, and it can be effectively identified by our token-wise contrastive excess score.
@@
We now proceed to explore the impact of token selection ratios ($k\%$) in our framework. Fig. 2 presents three representative curves: one task (BBH) under two transfer settings, and another task (News Headline Generation) under the same transfer setting for comparison. For BBH with the same backbone on source and target (Mistral 7B→Mistral 7B), a moderate token selection ratio of 30-70% yields the best results. Very low ratios (10-20%) lead to underfitting, while a 100% ratio is ineffective as it applies no filtering at all. Interestingly, in a weak-to-strong transfer (Llama2 7B → Llama3 8B) for the same BBH task, performance improves as k% decreases. This suggests that selecting only the tokens with the highest excess loss effectively filters out the majority of noisy regions where the weaker model is uncertain. By discarding the majority of tokens where the weaker model lacks confidence, TiTok can significantly reduce negative transfer. This trend, however, reverses for the News Headline Generation task under the same weak-to-strong transfer setting. For this task, a larger k% is generally better, potentially because even a weaker model can provide valuable lexical and stylistic cues that are useful for personalization. Consequently, unlike reasoning-focused tasks such as BBH, which are sensitive to noisy supervision, personalization tasks like News Headline Generation benefit more broadly from the source model’s outputs, even when the source model is relatively weaker than the target model.
Impact of query generation model.
We now examine how the choice of query generation model influences performance. Practically, synthetic queries can be generated either by the source model or by the target model, with the latter being the approach originally adopted in TransLoRA’s pipeline (Wang et al., 2024). Figure 3 compares these two options across different transfer settings, with the reported scores averaged over all BBH tasks. Notably, we observe that using the source expert model (source backbone + LoRA) for query generation generally yields stronger performance than using the target model. One possible explanation is that when synthetic data is generated by the same model, they remain closer to its training distribution. This alignment between queries and labels likely provides more coherent supervision, thereby facilitating more effective transfer. Ultimately, these findings empirically justify our design choice to use the source model as both the query and label generator, instead of using the target model.
@@
Table 4: Transfer using external data. ROUGE-1 (R-1) and ROUGE-L (R-L) on News Headline and Scholarly Title Generation under three data settings for Mistral 7B→Mistral 7B: (1) other-user, (2) mixed-user, (3) cross-task. Best scores are highlighted in bold.
@@
Table 5: Larger scale transfer settings. BBH evaluation results for cross-architecture and large-scale model transfers.
Transfer Setting Method Acc.
[…]
TiTok (k=70%) 0.463
Llama-3.3-70B $\rightarrow$ Llama-3.3-70B [^7] https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct Vanilla 0.593
TiTok (k=70%) 0.621
Transfer under cross-architecture and large-scale.
We additionally conduct experiments on larger-scale models and architectures to reinforce TiTok’s effectiveness. Due to the large model size, all experiments in this section are evaluated on BBH using 4-bit quantization. As shown in Table 5, TiTok successfully transfers token-level knowledge in both settings. Specifically, cross-architecture transfer from a dense model (Mistral 7B) to a mixture-of-experts (MoE) model yields a 1.98% improvement over the target baseline, while scaling up to the 70B transfer setting, achieves a 4.72% gain. These results confirm that TiTok generalizes robustly across both complex MoE architectures and massive model scales.
@@ 5 Conclusion
In this paper, we propose TiTok, a novel and efficient framework that transfers LoRA knowledge from a source model to a target model by training only on a selectively chosen set of highly informative tokens. At its core, TiTok strategically leverages token-level signals to distill task-relevant information from the source adapter, rather than indiscriminately relying on the entire token sequence. With this simple yet effective design, TiTok consistently surpasses baselines in various transfer settings, providing a practical solution for efficient knowledge transfer.
Limitations and Future Work
While TiTok has shown clear advantages, opportunities remain to refine the framework. First, generating synthetic data still requires a few seed examples. Nevertheless, this reliance is modest, as we avoid using full datasets. In addition, our experiments show external data serves as an effective alternative, reinforcing the flexibility of TiTok. Regarding token selection, while TiTok currently relies on a simple yet highly effective fixed threshold for token selection, future work could explore adaptive thresholding strategies to further enhance efficiency.
@@
We provide a comprehensive description of our implementation in Section 4, including pipeline configurations, hyperparameters, models, datasets, and evaluation metrics. The source code of TiTok is publicly available at https://github.com/NaughtyMaltiz16/TiTok.
Acknowledgments
This research was supported in part by Institute for Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2020-II201361, Artificial Intelligence Graduate School Program (Yonsei University); No. RS-2025-25442405, Development of a Self-Learning World Model-Based AGI System for Hyperspectral Imaging).
@@ Appendix A Tokenizer Alignment Algorithm
This process yields fractional mask scores that capture the relative importance of each target token. Finally, we keep only the top-$k\%$ of tokens according to these scores, producing a final binary selection mask over the target tokens that preserves the most informative regions while discarding less relevant ones. An overview of this tokenizer alignment algorithm is presented in Figure 4.
Robustness of the algorithm. The algorithm is conceptually error-free since it simply performs a deterministic mapping between two tokenizers on the exact same text sequence. Since both tokenizations correspond to the identical underlying character string, the mapping cannot introduce semantic errors; every token in both tokenizers is defined over non-overlapping spans of the same text, and these spans align uniquely. For instance, in the Mistral 7B → Llama3 8B transfer setting on BBH and News Headline benchmarks, 100% of tokens are aligned, and there are no exception cases as shown in Table 6, confirming that alignment errors do not occur in practice.
Table 6: Average alignment percentage breakdown across BBH and News Headline tasks in the Mistral 7B → Llama3 8B transfer setting.
BBH News Headline
exceptions 0% 0%
many to many 46.18% 5.00%
many to one 3.47% 17.45%
one to many 0.01% 0.01%
one to one 50.33% 77.52%
[…]
black
Furthermore, we additionally construct a degraded setting in which we keep only the one-to-one aligned token pairs and discard all other alignment cases. We perform the experiment in the Mistral 7B → Llama3 8B transfer setting on BBH and News Headline tasks, and the full results are presented in Table 7. From these empirical results, we can observe that using all the full alignment cases outperforms using one-to-one token pairs only. This implies that the many-to-one, one-to-many, and many-to-many alignments are also correctly aligned, as including them yields the best results. Together, these results confirm that our alignment method is both accurate and robust.
Table 7: Effect of token alignment cases in the Mistral-7B → Llama-3 8B transfer experiments.
BBH News Headline
Acc. R-1 R-L
Mistral-7B → Llama-3 8B Vanilla 0.397 0.117 0.101
One to one only 0.472 0.138 0.120
TiTok (k=70%) 0.482 0.160 0.142
[…]
black
Appendix B Usage of AI Assistants
AI assistants are minimally used in this work, restricted solely to language refinement, such as grammar correction, punctuation, and sentence structure. All research ideas, methodologies, and analyses are the original contributions of the authors. Thus, the use of AI is confined to editorial support and did not influence the originality or intellectual contributions of the work.
Appendix C Details of Datasets
C.1 Big-Bench Hard (BBH)
@@
C.2 Massive Multitask Language Understanding (MMLU)
Massive Multitask Language Understanding (MMLU) (Hendrycks et al., 2021) is a comprehensive benchmark for evaluating model performance across a broad range of knowledge intensive tasks. The benchmark consists of multiple choice questions and, similarly to BBH, we also split the 100 randomly sampled test-only instances into 90%/10%. All the 57 subtasks are categorized in Table 9.
@@ C.2 Massive Multitask Language Understanding (MMLU)
STEM abstract_algebra, anatomy, astronomy, college_biology, college_chemistry, college_computer_science, college_mathematics, college_medicine, college_physics, computer_security, conceptual_physics, electrical_engineering, elementary_mathematics, formal_logic, high_school_biology, high_school_chemistry, high_school_computer_science, high_school_mathematics, high_school_physics, high_school_statistics, machine_learning, medical_genetics, nutrition, professional_medicine, professional_psychology, virology, clinical_knowledge, human_aging, human_sexuality
Humanities business_ethics, formal_fallacies, jurisprudence, logical_fallacies, philosophy, prehistory, world_religions, moral_disputes, moral_scenarios, professional_law, professional_accounting, high_school_world_history, high_school_us_history, high_school_european_history, global_facts, security_studies
Social Sciences econometrics, high_school_macroeconomics, high_school_microeconomics, management, marketing, public_relations, sociology, us_foreign_policy, high_school_geography, high_school_government_and_politics
Other miscellaneous, global_facts
C.3 LaMP tasks
@@
  • News Headline Generation (LaMP 4). Given an author profile consisting of previously written headlines, the model is asked to generate a headline for a news article. The task evaluates if the model can adapt its output to reflect the author’s characteristic style in journalistic writing.
@@
Appendix D Baseline Details
@@ D.1 Vanilla
D.2 Knowledge Distillation (KD) (+MinED)
@@
D.3 TransLoRA
The TransLoRA baseline (Wang et al., 2024) transfers LoRA knowledge by generating synthetic data. In this method, the vanilla target model synthesizes queries, and the source model with its LoRA adapter provides the corresponding labels. Subsequently, a discriminator, separately trained with the source model as its base, is applied to filter the synthetic data used for fine-tuning the target model’s LoRA adapter. For the sake of consistency and fair comparison, we generate synthetic data both from the source model and use the same hyperparameter settings in our experiments .
Comparison with TransLoRA. TransLoRA’s core assumption is to transfer entire synthetic sequences generated by the teacher. Therefore, its framework is tightly coupled to synthetic data and is unable to function in the absence of teacher-generated labels. In that sense, this also implies that TransLoRA cannot be applied to external datasets. Thus, it mainly highlights the role of synthetic data, giving less thought to the architecture of the knowledge transfer procedure itself.
In contrast, TiTok directly rethinks the mechanism of knowledge transfer itself. TiTok concentrates on determining which tokens actually contain expert-specific knowledge (i.e., the tokens where the expert significantly deviates from the basic model) rather than copying entire synthetic sequences. Our novel token-level contrastive signal fundamentally changes how knowledge is extracted and aligned. Because this process does not rely on teacher-generated labels, TiTok naturally extends to external, non-synthetic datasets (See Table 4) and is therefore much more generalizable.
Therefore, despite the fact that the ultimate objective of both approaches is to solve the same high-level LoRA transfer problem, which inevitably leads to certain overlapping components, the underlying philosophies are essentially different. TransLoRA focuses on transferring full outputs and relies entirely on synthetic data, whereas TiTok focuses on transferring the expert’s internal knowledge signals regardless of the data source, thus being more generalizable and effective.